TRACTOR SKIDDING HYPOTHESIS TESTING Forest engineers are interested in studying the skidding distances of tractors along a new road in a European Forest. The engineers collect data on a random sample of 12 tractors. The collected data (in meters) are: 350, 285, 574, 439, 295, 184, 261, 273, 400, 311, 141, 425. The average skidding distance for the sample of 12 tractors is 328.17 and the standard deviation is 118.46. Local loggers working on the road believe that the mean skidding distance of all tractors on this road is 325 meters. The engineers believe it is much greater than this. Conduct a hypothesis test based on the engineer’s belief. Conduct this hypothesis test at a significance level of .01. What is the value of the calculated test statistic? Use the rounded standard error from question 1 and round the final answer to the thousandths place.
Answer #1

Sol:

Ho:μ 325

Ha:μ > 325

alpha=0.01

test statistic

t=xbar-mu/s/sqrt(n)

=(328.17 -325)/(118.46/sqrt(12)

t=0.093

the value of the calculated test statistic=t=0.093(rounded to 3 decimals)

degress of freedom=n-1=12-1=11

p value in excel is

=T.DIST.RT(0.093;11)

=0.463787893

=0..4639

p>0.01

Fail to reject Ho.

Accept Ho.

There is no sufficient statistical evidence at 1% level of significance to support the engineer's belief that

the mean skidding distance of all tractors on this road is greater than 325 meters.